A student titrates a 23.22 mL sample of HCl of unknown concentration. She begins her titration at the 0.00 mL mark and ends the titration at the 34.48 mL mark. The NaOH solution has a molarity of 1.057. What is the concentration of HCl?


1
Expert's answer
2021-03-29T06:04:10-0400

HCl-> NaCl + H20

mole ratio NaOH : HCl is 1:1

Molarity of Acid × V = Molarity of Base × V

= 34.48 - 23.22 = 11.26

23 × Molarity of HCl = 1.057 × 11.26

23 × M of HCl = 11.902

Concentration of HCl= 0.52